### v1

@@ Acknowledgements.
We thank our anonymous referee for their careful reading of our manuscript and valuable suggestions, which improved the quality of this work, particularly with regard to the realistic estimate of the gas volume density in the Oph-A cloud. We thank Bhaswati Mookerjea for the valuable comments and discussion at the early stage of the project and for kindly sharing the final data cube of the HCO+(4–3) transition line. We thank Fabio Santos for sharing the maps of dust temperatures and gas column density. We also thank Alex Lazarian for the constructive discussions concerning the DCF and DMA methods. N.L. thanks Thang Nguyen Tat for his valuable comments as part of the 3D structure of the Oph-A. N.L. and A.K. acknowledge support from the First TEAM grant of the Foundation for Polish Science No. POIR.04.04.00-00-5D21/18-00 (PI: A. Karska). A.K. acknowledges also support from the Polish National Agency for Academic Exchange grant No. BPN/BEK/2021/1/00319/DEC/1. N.B.N. was funded by the Master, Ph.D. Scholarship Programme of Vingroup Innovation Foundation (VINIF), code VINIF.2022.TS083. M.H. acknowledges support by the National Science Centre through the OPUS grant no. 2015/19/B/ST9/02959. N.L. also acknowledges support by the grant from the Simons Foundation to IFIRSE, ICISE (916424, PI: N.H.). P.N.D., N.B.N., and N.T.P. were funded by the Vietnam Academy of Science and Technology under project code THTETN.03/24-25. This work is based on observations made with the NASA/DLR Stratospheric Observatory for Infrared Astronomy (SOFIA). SOFIA is jointly operated by the Universities Space Research Association, Inc. (USRA), under NASA contract NNA17BF53C, and the Deutsches SOFIA Institut (DSI) under DLR contract 50 OK 2002 to the University of Stuttgart. This work made use of data from the JCMT/HARP and APEX. JCMT is operated by the East Asian Observatory on behalf of National Astronomical Observatory of Japan; the UK STFC under the auspices of grant number ST/N005856/1; Academia Sinica Institute of Astronomy and Astrophysics; the Korea Astronomy and Space Science Institute; the Operation, Maintenance and Upgrading Fund for Astronomical Telescopes and Facility Instruments, budgeted from the Ministry of Finance of China. This work used the facilities of the Canadian Astronomy Data Centre (CADC) operated by the National Research Council of Canada.References
@@ Appendix A Histogram of SOFIA/HAWC dataset
Figures 13 and 14 show the histogram of Stokes-, , SNRI, , , and SNRp of the SOFIA/HAWC data in band C and band D. Table 3 shows the summary of the mean and RMS of the distribution of Stokes-, , their uncertainties and SNRs.
@@ Appendix B Magnetic field orientations and strengths at different wavelengths
In Figure 15, we compare the B-fields vectors obtained from HAWC data at 89 and 154 m (this study), and from POL-2 data at 850 m (Kwon et al. 2018). Vectors show the mean B-field position angles at the subregions in Oph-A.
Figure 16 shows the maps of B-field strength estimated using the ST method () at 89 and 154 m. Figure 17 shows the histogram distribution of the obtained with HAWC+ data at 89 and 154 m.
@@ Appendix C Velocity dispersion traced by other molecular lines
To test whether the FWHM map of HCO+(43) line is a good choice, we further use the N2H+(32) molecular data observed by APEX to estimate the total velocity dispersion, . The majority of the Oph-A cloud exhibits the N2H+(32) emission showing 3 single Gaussian profiles; the middle peak is much stronger compared with the other two components. This strongest hyperfine line indicates the central velocity which is in the range of 2.54–3.97 km s-1, with an average velocity of 3.42 km s-1, in agreement with the systemic velocity of Oph-A of 3.44 km s-1 estimated by André et al. (2007). To each pixel we fit the spectrum with the sum of 3 single Gaussian profiles and then use the line width of the primary peak as the toward the computing pixel. We construct the map of the FWHM velocity dispersion, following the procedure presented in Section 3.2.2 for the case of the HCO+ molecular line.
@@
Figure 18 shows the map of using the N2H+(32) molecular line toward Oph-A. The values ranges from 0.18 to 1.73 km s-1 with a median value of 0.95 km s-1, in good agreement with those obtained by HCO+(43) line. We note, however, that the spatial extent of the HCO+(43) emission line is large than that of the N2H+(32). Thus, we use the map of FWHM of the HCO+(43) line to calculate the strengths in Oph-A. Similarly, White et al. (2015) reported a mean FWHM velocity dispersion of 1.5 km s-1 using the C18O(32) data from JCMT/HARP in overall entire Ophiuchus cloud. Using the N2H+(10) data from IRAM, André et al. (2007) estimated an average value of the 1D nonthermal velocity dispersion toward Oph-A, =0.2 km s-1, corresponding to 0.47 km s-1. Observations of the NH3 hyperfine emission from the Green Bank Telescope (GBT) also returned a similar 1D velocity dispersion with the average value of 0.3 km s-1, corresponding to an average of FWHM of 0.65 km s-1 (Chen et al. 2019).
@@
Appendix D Mass-to-flux ratio, Mach Alfvén Mach number, and beta plasma
@@
Figure 19 show the maps of Mass-to-flux ratio (top), Mach Alfvén Mach number (middle), and beta plasma (bottom) toward Oph-A using in band C. Figure 20 show maps of Mass-to-flux ratio (top), Mach Alfvén Mach number (middle), and beta plasma (bottom) toward Oph-A using in band C (left panels) and D (right panels).
Appendix E relation using estimated with the ST method
Figure 21 shows the relation between and where B-field strength is estimated using the St method.
Appendix F Main structure of the ridge identification
@@
We identify the elongated shape of the ridge in the densest region of Oph-A using the Python package RadFil. We first apply a threshold of 0.5 Jy/arcsec2 in the map of Stokes- at 154 m to build a mask for the main shape of the densest region (the ridge). The spine of the ridge is identified by using the Python package FilFinder777https://github.com/e-koch/FilFinder from Koch & Rosolowsky (2015). We then smooth the ridge spine and define a tangent line at every position on the ridge where the orientation of the ridge () at every position is identified. Figure 22 shows the schematic view of the ridge and the perpendicular cuts at each position on the ridge. The position angle of the tangent line at each position located on the ridge is defined as the orientation of the ridge at that position, measured east of north, and ranges from 0° to 180°.

### v2

@@ Acknowledgements.
We thank our anonymous referee for their careful reading of our manuscript and valuable suggestions, which improved the quality of this work, particularly with regard to the realistic estimate of the gas volume density in the Oph-A cloud. We thank Bhaswati Mookerjea for the valuable comments and discussion at the early stage of the project and for kindly sharing the final data cube of the HCO+(4–3) transition line. We thank Fabio Santos for sharing the maps of dust temperatures and gas column density. We also thank Alex Lazarian for the constructive discussions concerning the DCF and DMA methods. N.L. thanks Nguyen Tat Thang for his valuable comments as part of the 3D structure of the Oph-A. N.L. and A.K. acknowledge support from the First TEAM grant of the Foundation for Polish Science No. POIR.04.04.00-00-5D21/18-00 (PI: A. Karska). A.K. acknowledges also support from the Polish National Agency for Academic Exchange grant No. BPN/BEK/2021/1/00319/DEC/1. N.B.N. was funded by the Master, Ph.D. Scholarship Programme of Vingroup Innovation Foundation (VINIF), code VINIF.2023.TS.077. M.H. acknowledges support by the National Science Centre through the OPUS grant no. 2015/19/B/ST9/02959. N.L. also acknowledges support by the grant from the Simons Foundation to IFIRSE, ICISE (916424, PI: N.H.). P.N.D. and N.T.P. were funded by the Vietnam Academy of Science and Technology under project code THTETN.03/24-25. This work is based on observations made with the NASA/DLR Stratospheric Observatory for Infrared Astronomy (SOFIA). SOFIA is jointly operated by the Universities Space Research Association, Inc. (USRA), under NASA contract NNA17BF53C, and the Deutsches SOFIA Institut (DSI) under DLR contract 50 OK 2002 to the University of Stuttgart. This work made use of data from the JCMT/HARP and APEX. JCMT is operated by the East Asian Observatory on behalf of National Astronomical Observatory of Japan; the UK STFC under the auspices of grant number ST/N005856/1; Academia Sinica Institute of Astronomy and Astrophysics; the Korea Astronomy and Space Science Institute; the Operation, Maintenance and Upgrading Fund for Astronomical Telescopes and Facility Instruments, budgeted from the Ministry of Finance of China. This work used the facilities of the Canadian Astronomy Data Centre (CADC) operated by the National Research Council of Canada.References
@@
- Kandori et al. (2006) Kandori, R., Kusakabe, N., Tamura, M., et al. 2006, in Ground-based and Airborne Instrumentation for Astronomy, ed. I. S. McLean & M. Iye, Vol. 6269, International Society for Optics and Photonics (SPIE), 626951
@@ Appendix A Histogram of SOFIA/HAWC dataset
Figures 13 and 14 show histograms of Stokes-, , SNRI, , , and SNRp of the SOFIA/HAWC data in band C and band D. Table 3 shows the summary of the mean and RMS of the distribution of Stokes-, , their uncertainties and SNRs.
@@ Appendix B Magnetic field orientations and strengths at different wavelengths
In Fig. 15, we compare the B-fields vectors obtained from HAWC data at 89 and 154 m (this study), and from POL-2 data at 850 m (Kwon et al. 2018). Vectors show the mean B-field position angles at the subregions in Oph-A.
Figure 16 shows the histogram distribution of the estimated obtained with HAWC data at 89 and 154 m. Figure 17 shows the maps of B-field strength estimated using the ST method () at 89 and 154 m.
@@ Appendix C Velocity dispersion traced by other molecular lines
To test whether the FWHM map of HCO+(43) line is a good choice, we further use the N2H+(32) molecular data observed by APEX to estimate the total velocity dispersion, . The majority of the Oph-A cloud exhibits the N2H+(32) emission showing 3 single Gaussian profiles; the middle peak is much stronger compared with the other two components. This strongest hyperfine line indicates the central velocity which is in the range of 2.54–3.97 km s-1, with an average velocity of 3.42 km s-1, in agreement with the systemic velocity of Oph-A of 3.44 km s-1 estimated by André et al. (2007). To each pixel we fit the spectrum with the sum of 3 single Gaussian profiles and then use the line width of the primary peak as the toward the computing pixel. We construct the map of the FWHM velocity dispersion, following the procedure presented in Sect. 3.2.2 for the case of the HCO+(43) molecular line.
@@
Figure 18 shows the map of using the N2H+(32) molecular line toward Oph-A. The values ranges from 0.18 to 1.73 km s-1 with a median value of 0.95 km s-1, in good agreement with those obtained by HCO+(43) line. We note, however, that the spatial extent of the HCO+(43) emission line is largeer than that of the N2H+(32). Thus, we use the map of FWHM of the HCO+(43) line to calculate the strengths in Oph-A. Similarly, White et al. (2015) reported a mean FWHM velocity dispersion of 1.5 km s-1 using the C18O(32) data from JCMT/HARP in overall entire Ophiuchus cloud. Using the N2H+(10) data from IRAM, André et al. (2007) estimated an average value of the 1D nonthermal velocity dispersion toward Oph-A, =0.2 km s-1, corresponding to 0.47 km s-1. Observations of the NH3 hyperfine emission from the Green Bank Telescope also returned a similar 1D velocity dispersion with the average value of 0.3 km s-1, corresponding to an average of FWHM of 0.65 km s-1 (Chen et al. 2019).
@@
Appendix D Mass-to-flux ratio, Alfvénic Mach number, beta plasma, and relation results obtained by using
@@
Figure 19 shows the maps of mass-to-flux ratio (top panel), Alfvénic Mach number (middle panel), and beta plasma (bottom panel) toward Oph-A using in band C. Figure 20 show maps of mass-to-flux ratio (top), Alfvénic Mach number (middle), and beta plasma (bottom) toward Oph-A using in band C (left panels) and D (right panels).
Figure 21 shows the relation between and using in band D.
Appendix E Main structure of the ridge identification
@@
We identified the elongated shape of the ridge in the densest region of Oph-A using the Python package RadFil. We first applied a threshold of Stokes-0.5 Jy/arcsec2 in the map of Stokes- at 154 m to build a mask for the main shape of the densest region (the ridge). The spine of the ridge was identified by using the Python package FilFinder777https://github.com/e-koch/FilFinder from Koch & Rosolowsky (2015). We then smoothed the ridge spine and define a tangent line at every position on the ridge, where the orientation of the ridge () at every position was identified. Figure 22 shows the schematic view of the ridge and the perpendicular cuts at each position on the ridge. The position angle of the tangent line at each position located on the ridge is defined as the orientation of the ridge at that position, measured east of north, and ranges from 0° to 180°.